Using the ε-N definition of convergence of sequences, prove that limit of (2/n) when n goes to infinity is 0.
1
Expert's answer
2020-12-21T18:58:42-0500

We begin by examining the size of the difference:

"\big|\dfrac{2}{n}-0\big|=\dfrac{2}{n}"

Given "\varepsilon>0"


"\big|\dfrac{2}{n}-0\big|=\dfrac{2}{n}<\varepsilon, \text{if}\ n>\dfrac{2}{\varepsilon}"

Therefore using the limit definition


"\lim\limits_{n\to \infin}(\dfrac{2}{n})=0"

The sequence converges to  when "n" goes to infinity.